Determine whether the conditions for the binomial distribution are satisfied. If the conditions are not satisfied, state why not.

A polling agency in one city randomly selects 1000 registered voters from a pool of 385,600 registered voters of whom 17% are African American. Let X be the number of African Americans in the sample.

A.

The conditions for binomial distribution are not satisfied. The probability of success is not the same on each trial.

B.

The conditions for binomial distribution are not satisfied. There are more than two possible outcomes per trial.

C.

The conditions for binomial distribution are not satisfied. The trials are not independent. The sampling is without replacement and the sample size is less than 10% of population.

D.

The conditions for binomial distribution are satisfied.

Can you explain why the answer is what it is in detail, please.
Answer #1

option D is correct

The conditions for binomial distribution are satisfied.

(here below are the conditions for binomial distribution which are satisfied:

there are two outcomes : either African American or not African American

number of trails are fixed : n=1000

also as n/N =1000/385600 =0.0026 <=0.1 ; therefore we can consider the events to be independent)